# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ Abstract.
Famous results state that the classical MacWilliams identities fail for the Lee metric, the homogeneous metric and for the subfield metric, apart from some trivial cases. In this paper we change the classical idea of enumerating the codewords of the same weight and choose a finer way of partitioning the code that still contains all the information of the weight enumerator of the code. The considered decomposition allows for MacWilliams-type identities which hold for any additive weight over a finite chain ring. For the specific cases of the homogeneous and the subfield metric we then define a coarser partition for which the MacWilliams-type identities still hold. This result shows that one can, in fact, relate the code and the dual code in terms of their weights, even for these metrics. Finally, we derive Linear Programming bounds stemming from the MacWilliams-type identities presented.
Key words and phrases:
Lee weight, homogeneous weight, subfield weight, MacWilliams identity, Codes over rings
@@ 1. Introduction
The MacWilliams identities have been a major breakthrough in understanding the theory of linear codes and mark one of the most celebrated results. In a nutshell, the MacWilliams identities state that the weight enumerator of a code is related to, and completely determines, the weight enumerator of the dual code. The original results have been proven in [31] by Jessie MacWilliams and several generalizations and reformulations, through association schemes, generating functions or graphs, have appeared since [15, 27, 28, 42, 26]. Other generalizations of the MacWilliams identities have been discussed in [39] and further extensions to different code families, like convolutional codes, quantum codes and others, followed [11, 20, 21, 9, 46]. The MacWilliams identities also allow to state a Linear Programming (LP) bound, giving the tightest known upper bound on the size of a code with prescribed minimum distance and length. For an overview we refer the interested reader to [16].
While this powerful original result restricts itself to classical codes over finite fields endowed with the Hamming metric, many works have considered different ambient spaces and different metrics, with various outcomes. Some works on MacWilliams identities investigate partitions of the ambient space. Indeed, it is well-known that if the partition of the ambient space is Fourier-reflexive, the MacWilliams-type identities for this partition exist [19]. This last approach is mainly used for codes over rings and group codes, e.g. in [11, 25, 48]. The MacWilliams identities were generalized in many different ways, for instance, in [42] Wood employed a similar technique as described in [2] to generalize them to codes over Frobenius rings endowed with the Hamming metric. In 1972, Delsarte generalized them even further to additive codes using the Hamming metric [14].
However, as soon as we change the metric, the situation becomes more complex. While the MacWilliams identities are still valid for codes over $\mathbb{Z}/{4}\mathbb{Z}$ equipped with the Lee metric [24], in 2015, Shi et al. [37] and [43] proved their failure over $\mathbb{Z}/{8}\mathbb{Z}$. Moreover, Abdelghany and Wood [1] showed the nonexistence of any version of the MacWilliams identities for Lee weight enumerators over $\mathbb{Z}/{m}\mathbb{Z}$ for any positive integer $m\geq 5$. A similar result holds for the homogeneous weight as well: apart from the trivial cases $\mathbb{Z}/{p}\mathbb{Z}$, for a prime $p$, where the homogeneous metric coincides with the Hamming metric, the homogeneous MacWilliams identities fail [19, 44]. In both the cases of Lee and homogeneous metric, as well as for the subfield metric [22], one of the main problems when considering the weight distribution is the lack of knowledge about the entries of codewords for a fixed weight.
@@
Changing the partition of a code and, more precisely, moving away from enumerating codewords of the same weight, has already been suggested by MacWilliams herself for Lee-metric codes over finite fields [32]. Similarly, this idea has been supported for Lee-metric codes using association schemes in [6, 41] where the existence of MacWilliams-type identities have been shown. The study of partitioning the ambient space according to a pre-defined property dates back to Zinoviev and Ericson [47] in 1996. They studied additive codes and analyzed for which enumerators of the code and its dual MacWilliams-type identities are fulfilled, which laid the foundation of Fourier-reflexive partitions. In [19], Gluessing-Luerssen studied partitions of codes over Frobenius rings and showed that MacWilliams identities exist if the partition of the underlying ambient space is Fourier-reflexive. In a follow-up work [18], she proved that partitioning a code in the homogeneous metric into the codewords of the same homogeneous weight yields to a failure of the MacWilliams identities. Similarly, for codes in the sum-rank metric, [10] showed the failure of the classical MacWilliams identities and the existence of a MacWilliams-type identities using a finer partition.
We place this paper in this line of research and generalize the MacWilliams identities further. Instead of enumerating tuples (or vectors) of a given weight, we introduce a novel partition and show that it allows to state MacWilliams-type identities for any finite chain ring and any additive weight. Remarkably, the considered partition allows to retrieve the weight enumerator. While this result holds in general (as shown also in [19, 45]), we additionally put a greater focus on the three metrics for which the classical MacWilliams identities famously fail, namely the Lee metric, the homogeneous metric and the subfield metric. This not only shows that in this setting the weight enumerator of a code is related to (and completely determines) the weight enumerator of the dual, but also allows to state linear programming bounds for the three considered metrics.
@@
The paper is organized as follows. Section 2 will serve to introduce and recap the main definitions and results needed throughout the paper. We start with the notion of a linear code over a finite chain ring endowed with an additive weight (where we specifically introduce the Hamming, Lee, homogeneous and subfield weight). We then recap the theory of group characters stating some of their main properties, such as the Schur orthogonality. The section is completed by introducing the MacWilliams identities in their original form together with the Krawtchouk coefficients. Moving away from enumerating the codewords of the same weight, in Section 3 we present our novel results on the existence of MacWilliams-type identities. As a first step we consider a finite chain ring endowed with any additive metric and partition the code into codewords of the same (Lee weight) decomposition which we refer to as the Lee partition of the code. We show that the Lee decomposition enumerator of a code is fully determined by the Lee decomposition enumerator of its dual code, i.e., MacWilliams-type identities exist for any additive weight over any finite chain ring, and, in particular, for codes in the Lee metric. Even though the result holds for any additive metric over a finite chain ring, for some metrics there are coarser and more natural partitions to choose. Thus, in a next step we introduce partitions for the homogeneous weight and the subfield weight, respectively, that are more suited and coarser and still allow for MacWilliams-type identities. To complete the study, we derive linear programming bounds for the Lee, homogeneous and subfield metric in Section 4. Finally, concluding remarks are given in Section 5.
@@ 2. Preliminaries
In this section we introduce the main definitions and basic results needed throughout this paper. We denote by $p$ a prime number. Let $\mathcal{R}$ be a finite chain ring with maximal ideal generated by $\gamma$, let $s$ be its nilpotency index and $q=p^{r}$ be the size of the residue field, for a positive integer $r$. While the main result holds for any finite chain ring, we often focus on integer residue rings. We denote by $p$ a prime number and for a positive integer $s$ we let $\mathbb{Z}/{p^{s}}\mathbb{Z}$ denote the finite integer residue ring, and we denote by $(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$ the set of units modulo $p^{s}$. Note that a finite integer residue ring $\mathbb{Z}/{p^{s}}\mathbb{Z}$ is a finite chain ring with $q=p$, $\gamma=p$ and nilpotency index $s$.
Given a positive integer $n$ and an $\ell$-tuple of nonnegative integers $k:=(k_{1},\ldots,k_{\ell})$ satisfying that $\sum_{i=1}^{\ell}k_{i}=n$, we denote by
@@ 2.1. Codes and Additive Weights over Finite Chain Rings
A code $\mathcal{C}\subseteq\mathcal{R}^{n}$ can be represented by a generating set, that is a subset of codewords that generates $\mathcal{C}$ as an $\mathcal{R}$-submodule. We call a generating set a minimal generating set if it is minimal with respect to inclusion. The cardinality of a minimal generating set of $\mathcal{C}$ is called the rank $K$ and can be computed as $K=\sum_{i=0}^{s-1}k_{i}$. This allows for a matrix representation in terms of a minimal generating set.
Definition 2.
Given an $\mathcal{R}$-linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ of free rank $k_{0}$ and rank $K$, a matrix $G\in\mathcal{R}^{K\times n}$ whose rows form a generating set of $\mathcal{C}$ is called a generator matrix of the code. A matrix $H\in\mathcal{R}^{(n-k_{0})\times n}$ whose kernel coincides with $\mathcal{C}$ is called a parity-check matrix of $\mathcal{C}$.
Usually, it is convenient to consider a generator and parity-check matrix in their standard form defined in [36].
Proposition 3 ([36, Proposition 3.2]).
Let $\mathcal{R}$ be a finite chain ring and let $\mathcal{C}\subseteq\mathcal{R}^{n}$ be a linear code of subtype $(k_{0},\dots,k_{s-1})$. Then $\mathcal{C}$ is permutation equivalent to a code having a generator matrix in the following standard form
(4) $G=\begin{bmatrix}I_{k_{0}}&A_{0,1}&A_{0,2}&A_{0,3}&\dots&A_{0,s-1}&A_{0,s}\\ 0&\gamma I_{k_{1}}&\gamma A_{1,2}&\gamma A_{1,3}&\dots&\gamma A_{1,s-1}&\gamma A_{1,s}\\ 0&0&\gamma^{2}I_{k_{2}}&\gamma^{2}A_{2,3}&\dots&\gamma^{2}A_{2,s-1}&\gamma^{2}A_{2,s}\\ \vdots&\vdots&\vdots&\vdots&&\vdots&\vdots\\ 0&0&0&0&\dots&\gamma^{s-1}I_{k_{s-1}}&\gamma^{s-1}A_{s-1,s}\\ \end{bmatrix}\ ,$
where $A_{i,s}\in(\mathcal{R}/\gamma^{s-i}\mathcal{R})^{k_{i}\times(n-K)}$ and $A_{i,j}\in(\mathcal{R}/\gamma^{s-i}\mathcal{R})^{k_{i}\times k_{j}}$ for $j<s$. Moreover $\mathcal{C}$ is permutation equivalent to a code having a parity-check matrix in standard form
(5) $H=\begin{bmatrix}B_{0,0}&B_{0,1}&\dots&B_{0,s-1}&I_{n-K}\\ \gamma B_{1,0}&\gamma B_{1,1}&\dots&\gamma I_{k_{s-1}}&0\\ \vdots&\vdots&&\vdots&\vdots\\ \gamma^{s-1}B_{s-1,0}&\gamma^{s-1}I_{k_{1}}&\dots&0&0\\ \end{bmatrix},$
where $B_{0,j}\in(\mathcal{R}/\gamma^{s}\mathcal{R})^{(n-K)\times k_{j}},B_{i,j}\in(\mathcal{R}/\gamma^{s-i}\mathcal{R})^{k_{s-i}\times k_{j}},$ for $j>1.$
@@
It is easy to see that such an additive distance over $\mathcal{R}$ satisfies the properties of a metric. We will now introduce the main additive weight functions which are of interest in the course of this paper. Each of the weights over $\mathcal{R}$ induce a distance function as described above.
The most prominent weight in coding theory is the Hamming weight.
@@
(9) $\operatorname{wt}_{\mathsf{H}}(x):=\left|\,{\left\{{i=1,\ldots,n\;|\;x_{i}\neq 0}\right\}}\,\right|.$
We now focus on the case where $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$ is a finite integer residue ring modulo $p^{s}$. Hence, we consider codes to be submodules of $(\mathbb{Z}/p^{s}\mathbb{Z})^{n}$. In this work, we are particularly interested in the Lee weight and homogeneous weight. The Lee weight, introduced in 1958 by Lee [30], is a generalization of the Hamming weight over the binary.
@@
In 1997, Constantinescu and Heise first introduced the homogeneous weight for integer residue rings [13], which was later extended to finite rings in [23] and [35]. For this we consider again a finite chain ring $\mathcal{R}$ with socle $\mathcal{S}.$ Recall that the socle $\mathcal{S}$ of a ring is defined as the sum of the minimal nonzero submodules. For example, for $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$ we have $\mathcal{S}=\langle p^{s-1}\rangle$. More generally, if the maximal ideal of $\mathcal{R}$ is $\langle\gamma\rangle$, we have $\mathcal{S}=\langle\gamma^{s-1}\rangle.$
The homogeneous weight extends the Hamming weight coinciding with it for finite fields and with the Lee weight for $\mathbb{Z}/{4}\mathbb{Z}$. In the following we define the homogeneous weight in its normalized form.
@@
For the third metric, let us consider finite fields, more precisely finite extension fields $\mathbb{F}_{p^{r}}$. In [22], the authors introduced the subfield metric over $\mathbb{F}_{p^{r}}$ for quantum error correcting codes. This metric is defined similarly to the homogeneous metric over finite chain rings. Namely, it gives weight $1$ to nonzero elements in the underlying base field and all other nonzero elements in $\mathbb{F}_{p^{r}}$ are of weight $\lambda\geq 1$.
@@ 2.2. Group Characters
We now recall the required basics on group characters. In a nutshell, a group character defines a representation of a group $G$ in terms of a complex function. Working over cyclic groups, characters are used in combination with the discrete Fourier transform.
@@
Let $G$ be a finite abelian group. A character $\chi$ of $G$ is a complex-valued map $\chi:G\longrightarrow\left\{{z\in\mathbb{C}\;|\;\left|\,{z}\,\right|=1}\right\}$, such that for every $\alpha,\beta\in G$ it holds
@@
A character is called principal if it maps every element of the group to $1$.
@@
In the following, we derive explicit expressions for characters over certain finite abelian groups.
  1. (1)
    If $G=(\mathbb{Z}/{m}\mathbb{Z},+)$ with $m$ a positive integer, for any $a,x\in\mathbb{Z}/{m}\mathbb{Z}$ the characters of $G$ are given by
    (18) $\chi_{a}(x):=\xi^{ax}\ ,$
    where $\xi=e^{\frac{2\pi i}{m}}$ is an $m$-th root of unity.
  2. (2)
    Let $G$ be a finite abelian group. From the fundamental theorem of finite abelian groups, $G$ decomposes uniquely (up to the order of summands) into a finite direct sum of cyclic subgroups of prime power order, namely,
    (19) $G=\mathbb{Z}/{q_{1}}\mathbb{Z}\times\dots\times\mathbb{Z}/{q_{t}}\mathbb{Z}\ .$
    For any $a=(a_{1},\dots,a_{t}),\ x=(x_{1},\dots,x_{t})\in G$ any character has the form
    (20) $\chi_{a}(x)=\prod_{i=1}^{t}\xi_{i}^{a_{i}x_{i}}\ ,$
    where $\xi_{i}$ is a $q_{i}$-th root of unity. In particular, if $G=(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and $a,x\in G$ the characters are given by
    (21) $\chi_{a}(x)=\xi^{\left\langle{a,x}\right\rangle}\ ,$
    where $\left\langle{a,x}\right\rangle:=ax^{\top}$ and $\xi$ is a $p^{s}$-root of unity.
  3. (3)
    Let $G=(\mathbb{F}_{p^{r}}^{n},+)$ and $f:\mathbb{F}_{p^{r}}\longrightarrow\mathbb{F}_{p}$ a nontrivial $\mathbb{F}_{p}$-linear map. Any character of $G$ is given by $\chi_{a}(x)=\xi^{f(\left\langle{x,a}\right\rangle)}$, where $\xi$ is a primitive $p$-th root of unity and for $a,x\in G$, $\left\langle{a,x}\right\rangle=ax^{\top}$ denotes the standard inner product in $\mathbb{F}_{p^{r}}^{n}$. As an example of such non-trivial linear mapping, we can consider the expansion of an element $a\in\mathbb{F}_{p^{r}}$ to a vector $\Gamma(a)\in\mathbb{F}_{p}^{r}$ using a basis $\Gamma$ of $\mathbb{F}_{p^{r}}$ over $\mathbb{F}_{p}$ and project on the first coordinate of $\Gamma(a).$ Another example is the trace function $\operatorname{Tr}(x)=\sum_{i=0}^{m-1}x^{p^{i}},$ which we will use throughout this paper as $f$. Therefore, any character of $G$ has the form
    (22) $\chi_{a}(x)=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}\ ,$
    where $\xi$ is a primitive $p$-th root of unity.
  4. (4)
    Let $\mathcal{R}=\mathrm{GR}(p^{s},r)$ be the Galois ring of characteristic $p^{s}$ and with $p^{sr}$ elements, namely $\mathcal{R}$ is the Galois extension of $\mathbb{Z}/{p^{s}}\mathbb{Z}$ of degree $r$. Recall that any element $a\in\mathcal{R}$ has a unique $p$-adic representation, that is $a=a_{0}+a_{1}p+\dots+a_{s-1}p^{s-1}$, where, for any $11\leq i\leq s-$ 1, $a_{i}$ belongs to the Teichmüller set of $\mathcal{R}$ [33]. Over Galois rings, one can define the generalized Frobenius map as
    (23) $f\colon\mathcal{R}\to\mathcal{R},\quad a\mapsto a^{f}\coloneqq a_{0}^{p}+pa_{1}^{p}+\dots+p^{s-1}a_{s-1}^{p}\ .$
    Note that, if $r=1$, $f$ reduces to the Frobenius automorphism. The generalized Frobenius induces the generalized trace map
    (24) $\operatorname{Tr}\colon\mathcal{R}\to\mathbb{Z}/{p^{s}}\mathbb{Z},\quad a\mapsto a+a^{f}+a^{f^{2}}+\dots+a^{f^{r-1}}\ .$
    The generalized trace shares some properties with the trace map; for instance, the generalized trace is $\mathbb{Z}/{p^{s}}\mathbb{Z}$-linear, and it takes each value of $\mathbb{Z}/{p^{s}}\mathbb{Z}$ with equal frequency. For further details on this topic, refer to [40]. Similarly to the finite fields case, any character over $\mathcal{R}$ is of the form
    (25) $\chi_{a}(x)=\xi^{\operatorname{Tr}(ax)}\ ,$
    where $\xi$ is a $p^{s}$-th root of unity. Note that this generalizes the case of $\mathbb{Z}/{p^{s}}\mathbb{Z}$ when $r=1$, where the trace function coincides with the identity, as well as the case of $\mathbb{F}_{p^{r}}$ where $s=1$. In analogy with the finite abelian group case, characters in $\mathcal{R}^{n}$ can be explicitly described. Specifically, for any $a,x\in\mathcal{R}^{n}$ every character can be expressed as
    (26) $\chi_{a}(x)=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}\ ,$
    where $\xi$ is a $p^{s}$-th root of unity.
  5. (5)
@@
    As in the previous cases, we can derive expressions for characters in $\mathcal{R}^{n}$. Given $a=(a^{(1)},\ldots,a^{(n)}),x=(x^{(1)},\ldots,x^{(n)})\in\mathcal{R}^{n},$ where $a^{(i)}=(a^{(i)}_{1},\ldots,a^{(i)}_{t})\in\mathcal{R}$ all the characters are of the form
@@
Let $G$ be a finite group and $\chi$ a character of $G$. Then
@@
A similar result exists for subgroups (or linear codes). However, in that case, we distinguish between characters whose kernel is contained in the given subgroup and those whose kernel is not (see for instance [42]). More formally, let $H\subset G$ be a subgroup of a finite abelian group $G$. We denote by $\widehat{G}$ the set of all characters in $G$. Define the annihilator of $H\subset G$ as
@@
A character $\chi\in(\widehat{G}:H)$ is called a generating character. Then, the orthogonality relation in Lemma 12 extends to a subgroup as follows.
@@
(32) $\sum_{h\in H}\chi(h)=\begin{cases}\left|\,{H}\,\right|&\text{ if }\chi\text{ is a generating character,}\\ 0&\text{ otherwise.}\end{cases}$
@@
Let $G$ be a finite abelian group and let $H$ be a subgroup of $G$. For a non-generating character $\chi$ (i.e., $\chi\not\in(\widehat{G}:H)$) it holds that
@@
In particular, if $\mathcal{R}$ is a finite chain ring and $I$ is an ideal in $\mathcal{R}$, for a non-generating character $\chi$ over $\mathcal{R}$ we have $\sum_{x\in I}\chi(x)=0$, while $\sum_{x\in I\setminus\{0\}}\chi(x)=-1.$
We now turn our attention to partitions of abelian groups. Following [48], a partition $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ of $G$ consists of non-empty pairwise disjoint sets $P_{i}$, $1\leq i\leq m$, that cover $G$. The sets of a given partition are called blocks.

### v3

# Coarsest Fourier-reflexive Partitions for the Lee, Homogeneous and Subfield Metric

@@ Abstract.
MacWilliams identities relate the weight enumerators of a code with those of its dual and are classically formulated with respect to the Hamming weight. For other metrics, however, these identities often fail when considering the weight partition of the ambient space. It is known that MacWilliams identities hold for enumerators associated with Fourier-reflexive partitions, and that orbits of subgroups of the linear isometry group always yield such partitions. This raises the question whether, for metrics beyond the Hamming metric, there exist meaningful partitions that lie strictly between the weight partition and the orbit partition: finer than the latter, yet still coarse enough to retain useful MacWilliams-type identities. In this work, we study this question for finite chain rings endowed with additive metrics. For the Lee metric, we show that the partition induced by the action of the full group of linear isometries is already the coarsest Fourier-reflexive partition refining the weight partition. In particular, no intermediate partition exists that is both finer than the Lee weight partition and Fourier-reflexive. We refer to this partition as the Lee partition and show that it allows the recovery of all additive weight enumerators over the ring. In contrast, for the homogeneous metric and for the subfield metric, we identify new, significantly coarser symmetrized partitions that remain Fourier-reflexive and still allow the recovery of the corresponding weight enumerators. We prove that these partitions are the coarsest such symmetrized partitions for which MacWilliams-type identities hold. As an application, we derive linear programming bounds based on the resulting MacWilliams identities.
@@ 1. Introduction
The MacWilliams identities have been a major breakthrough in understanding the theory of linear codes and mark one of the most celebrated results. In a nutshell, the MacWilliams identities state that the weight enumerator of a code is related to, and completely determines, the weight enumerator of the dual code. The original result has been proven in [1] by Jessie MacWilliams and several generalizations and reformulations, through association schemes, generating functions or graphs, have appeared since [2, 3, 4, 5, 6]. Other generalizations of the MacWilliams identities have been discussed in [7] and further extensions to different code families, like convolutional codes, quantum codes and others, followed [8, 9, 10, 11, 12]. The MacWilliams identities also allow to state a Linear Programming (LP) bound, giving the tightest known upper bound on the size of a code with prescribed minimum distance and length. For an overview we refer the interested reader to [13].
@@
This powerful original result restricts itself to classical codes over finite fields endowed with the Hamming metric. However, as soon as we change the metric, the situation becomes more complex. While the MacWilliams identities are still valid for codes over $\mathbb{Z}/{4}\mathbb{Z}$ equipped with the Lee metric [14], in 2015, Shi et al. [15] and [16] proved their failure over $\mathbb{Z}/{8}\mathbb{Z}$. Moreover, Abdelghany and Wood [17] showed the nonexistence of any version of the MacWilliams identities for Lee weight enumerators over $\mathbb{Z}/{m}\mathbb{Z}$ for any positive integer $m\geq 5$. A similar result holds for the homogeneous weight as well: apart from the trivial cases $\mathbb{Z}/{p}\mathbb{Z}$, for a prime $p$, where the homogeneous metric coincides with the Hamming metric, the homogeneous MacWilliams identities fail [18, 19]. In [20, 21], the rarity of weights that satisfy the MacWilliams identities is observed. Changing the partition of a code and, more precisely, moving away from enumerating codewords of the same weight, has already been suggested by MacWilliams herself for Lee-metric codes over finite fields [22]. Similarly, this idea has been supported for Lee-metric codes using association schemes in [23, 24] where the existence of MacWilliams-type identities have been shown. The study of partitioning the ambient space according to a pre-defined property dates back to Zinoviev and Ericson [25] in 1996. They studied additive codes and analyzed for which enumerators of the code and its dual MacWilliams-type identities are fulfilled, which laid the foundation of Fourier-reflexive partitions. In [18], Gluessing-Luerssen, studied partitions of additive codes over finite Frobenius rings and showed that MacWilliams identities exist if the partition of the underlying ambient space is Fourier-reflexive. In a follow-up work [26], she proved that partitioning a code in the homogeneous metric into the codewords of the same homogeneous weight yields to a failure of the MacWilliams identities. Similarly, for codes in the sum-rank metric, [27] showed the failure of the classical MacWilliams identities and the existence of MacWilliams-type identities using a finer partition.
@@
In this paper, we focus on codes over a finite chain ring $\mathcal{R}$ endowed with an additive weight (i.e., the weight of a codeword is given by the sum of the weights of its entries over $\mathcal{R}$). For additive weights on codes $\mathcal{C}\subseteq\mathcal{R}^{n}$, it is natural to consider partitions of the ambient space $\mathcal{R}$ that induce symmetrized partitions of $\mathcal{R}^{n}$. MacWilliams-type identities for such symmetrized partition enumerators were studied in [18], where it was shown that they hold whenever the underlying partition of $\mathcal{R}$ is Fourier-reflexive. Classical examples include the Hamming weight partition, the singleton partition, and partitions induced by the orbits of subgroups of the additive automorphism group. These examples illustrate a fundamental trade-off. The Hamming weight partition is very coarse and satisfies MacWilliams identities, but it does not allow to recover other additive weight enumerators. At the opposite extreme, the singleton partition is Fourier-reflexive and allows the recovery of all additive weights, but it is maximally fine and becomes computationally infeasible as the alphabet size grows. Orbit partitions arising from isometry groups lie between these two extremes and are guaranteed to be Fourier-reflexive, but they are often still unnecessarily fine. The central goal of this work is therefore to identify, for a given additive metric, the coarsest symmetrized partition that is Fourier-reflexive and still determines the corresponding weight enumerator. We address this problem for finite chain rings endowed with the Lee metric, the homogeneous metric, and the subfield metric. In particular, we show that for the Lee metric no such intermediate partition exists: the orbit partition induced by the full isometry group is already the coarsest Fourier-reflexive refinement. In contrast, for the homogeneous and subfield metric we construct new, significantly coarser symmetrized partitions that retain MacWilliams-type identities while still allowing the recovery of the respective weight enumerators. The paper is organized as follows. Section 2 will serve as introduction and recap of the main definitions and results needed throughout the paper. We start with the notion of a linear code over a finite chain ring endowed with an additive weight (where we specifically introduce the Hamming, Lee, homogeneous and subfield weight). We then recap the theory of group characters stating some of their main properties, such as the Schur orthogonality. The section is completed by introducing the MacWilliams identities in their original form together with the Krawtchouk coefficients. Instead of partitioning tuples by weight, in Section 3 we present novel Fourier-reflexive partitions with respect to the Lee, homogeneous and subfield weight. As a first step we consider a finite chain ring endowed with any additive metric and partition the code into codewords of the same (Lee weight) decomposition which we refer to as the Lee partition of the code. We show that the Lee decomposition enumerator of a code is fully determined by the Lee decomposition enumerator of its dual code, i.e., MacWilliams-type identities exist for any additive weight over any finite chain ring, and, in particular, for codes in the Lee metric. Even though the result holds for any additive metric over a finite chain ring, for some metrics there are coarser and more natural partitions to choose. Thus, in a next step we introduce partitions for the homogeneous weight and the subfield weight, respectively, that are more suited and coarser and still allow for MacWilliams-type identities. To complete the study, we derive linear programming bounds for the Lee, homogeneous and subfield metric in Section 4. Finally, concluding remarks are given in Section 5.
@@ 2. Preliminaries
In this section we introduce the main definitions and basic results needed throughout this paper. We denote by $p$ a prime number. Let $\mathcal{R}$ be a finite chain ring with maximal ideal generated by $\gamma$, let $s$ be its nilpotency index and $q=p^{r}$ be the size of the residue field, for a positive integer $r$. While the main result holds for any finite chain ring, we often focus on integer residue rings. We denote by $p$ a prime number and for a positive integer $s$ we let $\mathbb{Z}/{p^{s}}\mathbb{Z}$ denote the finite integer residue ring, and we denote by $(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$ the set of units modulo $p^{s}$. Note that a finite integer residue ring $\mathbb{Z}/{p^{s}}\mathbb{Z}$ is a finite chain ring with $q=p$, $\gamma=p,r=1$ and nilpotency index $s$. Given a positive integer $n$ and an $\ell$-tuple of nonnegative integers $k:=(k_{1},\ldots,k_{\ell})$ satisfying that $\sum_{i=1}^{\ell}k_{i}=n$, we denote by
@@ 2.1. Codes and Additive Weights over Finite Chain Rings
A code $\mathcal{C}\subseteq\mathcal{R}^{n}$ can be represented by a generating set, that is a subset of codewords that generates $\mathcal{C}$ as an $\mathcal{R}$-submodule. We call a generating set a minimal generating set if it is minimal with respect to inclusion. The cardinality of a minimal generating set of $\mathcal{C}$ is called the rank $K$ and can be computed as $K=\sum_{i=0}^{s-1}k_{i}$.
@@
It is easy to see that such an additive distance over $\mathcal{R}$ satisfies the properties of a metric. We will now introduce the main additive weight functions which are of interest in the course of this paper. Each of the weights over $\mathcal{R}$ induce a distance function as described above. The most prominent weight in coding theory is the Hamming weight.
@@
(8) $\operatorname{wt}_{\mathsf{H}}(x):=\left|\,{\left\{{i\in\{1,\ldots,n\}\;|\;x_{i}\neq 0}\right\}}\,\right|.$
We now focus on the case where $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$ is a finite integer residue ring modulo $p^{s}$. Hence, we consider codes to be submodules of $(\mathbb{Z}/p^{s}\mathbb{Z})^{n}$. The Lee weight, introduced in 1958 by Lee [28], is a generalization of the Hamming weight over the binary.
@@
In 1997, Constantinescu and Heise first introduced the homogeneous weight for integer residue rings [29], which was later extended to finite rings in [30] and [31]. For this we consider again a finite chain ring $\mathcal{R}$ with socle $\mathcal{S}.$ Recall that the socle $\mathcal{S}$ of a ring is defined as the sum of the minimal nonzero submodules. For example, for $\mathcal{R}=\mathbb{Z}/{p^{s}}\mathbb{Z}$ we have $\mathcal{S}=\langle p^{s-1}\rangle$. More generally, if the maximal ideal of $\mathcal{R}$ is $\langle\gamma\rangle$, we have $\mathcal{S}=\langle\gamma^{s-1}\rangle.$ The homogeneous weight extends the Hamming weight coinciding with it for finite fields and with the Lee weight for $\mathbb{Z}/{4}\mathbb{Z}$. In the following we define the homogeneous weight in its normalized form.
@@
Lastly, let us consider finite fields, more precisely finite extension fields $\mathbb{F}_{p^{r}}$. In [32], the authors introduced the subfield metric over $\mathbb{F}_{p^{r}}$ for quantum error correcting codes. This metric is defined similarly to the homogeneous metric over finite chain rings. Namely, it gives weight one to nonzero elements in the underlying base field and all other nonzero elements in $\mathbb{F}_{p^{r}}$ are of weight $\lambda\geq 1$.
@@ 2.2. Group Characters
We now recall the required basics on group characters. In a nutshell, a group character defines a representation of a group $G$ in terms of a complex function. Working over cyclic groups, characters are used in combination with the discrete Fourier transform. Throughout this section, let $\mathbb{C}^{*}\coloneqq\left\{{z\in\mathbb{C}\;|\;\left|\,{z}\,\right|=1}\right\}$.
@@
Let $G$ be a finite abelian group. A character $\chi$ of $G$ is a complex-valued map $\chi:G\longrightarrow\mathbb{C}^{*}$, such that for every $\alpha,\beta\in G$ it holds
@@
We denote by $\widehat{G}$ the set of all characters of $G$, where $\widehat{G}=\operatorname{Hom}_{\mathbb{Z}}(G,\mathbb{C}^{*})$ forms a group along with the addition
(17) $(\chi_{i}+\chi_{j})(g)=\chi_{i}(g)\chi_{j}(g),\quad\text{for all }\chi_{i},\chi_{j}\in\widehat{G}\text{ and }g\in G.$
[…]
@@
We can derive expressions for characters in $\mathcal{R}^{n}$. Given $a=(a^{(1)},\ldots,a^{(n)}),x=(x^{(1)},\ldots,x^{(n)})\in\mathcal{R}^{n},$ where $a^{(i)}=(a^{(i)}_{1},\ldots,a^{(i)}_{t})\in\mathcal{R}$ all the characters are of the form
@@
Let $G$ be a finite group and $\chi$ a character of $G$. Then it holds that
@@
A similar result exists for subgroups (or linear codes). However, in that case, we distinguish between characters whose kernel is contained in the given subgroup and those whose kernel is not (see for instance [5]). More formally, let $H\subset G$ be a subgroup of a finite abelian group $G$. Define the annihilator of $H\subset G$ as
@@
Definition 11.
Let $\chi$ be a character of $G$ and let $H$ be a subgroup of $G$. Then $\chi$ is said to be generating for $H$, or simply generating, if $\chi\not\in(\widehat{G}:H)$, that is, $H\not\subset\ker(\chi)$.
The orthogonality relation in Lemma 10 extends to a subgroup as follows.
@@
(23) $\sum_{h\in H}\chi(h)=\begin{cases}0&\text{ if }\chi\text{ is generating for $H$,}\\ \left|\,{H}\,\right|&\text{ otherwise.}\end{cases}$
@@
Let $G$ be a finite abelian group and let $H$ be a subgroup of $G$. If $\chi$ is a generating character for $H$ (i.e., $\chi\not\in(\widehat{G}:H)$) it holds that
@@
In particular, if $\mathcal{R}$ is a finite chain ring and $I$ is an ideal in $\mathcal{R}$, for a generating character $\chi$ over $\mathcal{R}$ we have $\sum_{x\in I}\chi(x)=0$, while $\sum_{x\in I\setminus\{0\}}\chi(x)=-1.$
2.3. Partitions and Induced Partitions
We now turn our attention to partitions of a finite abelian group $G$. Following [35], a partition $\mathcal{P}=P_{1}\mid P_{2}\mid\dots\mid P_{m}$ of $G$ consists of non-empty pairwise disjoint sets $P_{i}$, $1\leq i\leq m$, that cover $G$. The sets of a given partition are called blocks.